# Asymptotics for partitions over the Fibonacci numbers and related sequences

### 12 Dec 2023 v1

@@ 5. Partitions over general linear recurrences with a dominant root
The results of the previous sections on Fibonacci partitions can be generalised to positive recurrence sequences $P_{n}$ with an irreducible characteristic polynomial having a positive dominant real root and such that $P_{1}=1$. We will also assume that the values $P_{n}$ are distinct. Here, we require $P_{1}=1$ so that there always exists a non-distinct partition of $n$ over $P_{n}$. The restriction on the values of $P_{n}$ being distinct is not a very strict one—since the recursion has a dominant real root there are at most finitely many repeated values, so an analysis analogous to moving between $F(x)$ and $F_{2}(x)$ (or, $G(x)$ and $G_{2}(x)$) is possible, and uncomplicated. For such $P_{n}$, we wish to asymptotically understand the number of solutions to
@@
As before, we let $p_{P}(n)$ denote the number of non-distinct partitions of $n$ over the sequence $P_{n}$, that is, solutions to (1) in nonnegative integers $a_{k}$, and let $q_{P}(n)$ denote the number of distinct partitions of $n$ over the sequence $P_{n}$, that is, solutions to (1) with $a_{k}\in\{0,1\}$.
The case of partitions over $P_{n}$ is carried out exactly as in the case with the Fibonacci numbers $F_{n}$, but with the Fibonacci zeta function replaced by the zeta function $\zeta_{P}(z)$ which is the meromorphic continuation of the Dirichlet series $\sum_{k\geqslant 1}P_{k}^{-z}$ ($\Re(z)>0$). Here, we consider the generating functions,
$F_{P}(x):=\sum_{n\geqslant 0}p_{P}(n)x^{n}=\prod_{k\geqslant 1}\left(1-x^{P_{k}}\right)^{-1}$
and
$G_{P}(x):=\sum_{n\geqslant 0}\left(\sum_{m\leqslant n}q_{P}(m)\right)x^{n}=(1-x)^{-1}\prod_{k\geqslant 1}\left(1+x^{P_{k}}\right).$
@@
  • a triple pole (resp., double pole) at $z=0$ in the case of $F_{P}(e^{-s})$ (resp., $G_{P}(e^{-s})$),
@@
Now, it is clear from the previous analysis on Fibonacci partitions that the main contribution will come from the pole at $z=0$, which comes only from $\boldsymbol{k}=\boldsymbol{0}=(0,\ldots,0).$ The negative real poles will give a cumulative contribution of $O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}})$ for any fixed $\varepsilon>0$, since $s\log s=O(s^{1-\varepsilon})$ for any fixed $\varepsilon>0$ as $s\to 0^{+}$. As well, the simple poles off the real line contribute a function $f_{2}(s)$ (resp., $f_{3}(s)$) toward $\log F_{P}(e^{-s})$ (resp., $\log G_{P}(e^{-s})$), that satisfy $f_{i}(s)=f_{i}(\beta s)+O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}})$ for $i\in\{2,3\}$. It remains to obtain the contributions from the pole at $z=0$. For these, we require the following result.
@@
In the following result, we use Lemma 4 to determine the asymptotic behaviour of the functions $\log F_{P}(e^{-s})$ and $\log G_{P}(e^{-s})$ as $s\to 0^{+}$. All of the expansions of the functions involved, $s^{-z}$, $\Gamma(z)$, $\zeta(1+z)$ and $\eta(1+z)$, have been noted somewhere in the previous sections of this work—we use them below without further reference.
@@
For any fixed $\varepsilon\in(0,1)$, as $s\to 0^{+}$, the functions $F_{P}(z)$ and $G_{P}(z)$ defined above satisfy,
@@
and
$\log G_{P}(e^{-s})=-\left(\frac{\log 2}{\log\beta}+1\right)\log s+C_{4}+f_{3}(s)+O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}),$
@@
$C_{4}:=2\gamma+C_{3}=2\gamma-\log 2\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)+\frac{1}{\log\beta}\left(2\gamma-\frac{(\log 2)^{2}}{2}\right),$
and $f_{i}(s)=f_{i}(\beta s)+O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}})$ for $i\in\{2,3\}$.
@@
For $F_{P}(e^{-s})$, note that near $z=0$, we have
@@
where $C_{2}:=\frac{1}{\log\beta}\left(\frac{1}{2}\left(\gamma^{2}+\frac{\pi^{2}}{6}\right)+\gamma^{2}-\gamma_{1}\right)-\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)2\gamma+\left(\frac{(\log\lambda)^{2}}{2\log\beta}+\frac{\log\lambda}{2}+\frac{\log\beta}{12}+C_{1}\right).$ The contributions coming from the rest of the poles are as discussed before the statement of Lemma 4.
Again, near $z=0$, considering now the case $G_{P}(e^{-s})$, we have
$s^{-z}\Gamma(z)\eta(1+z)=\frac{\log 2}{z}+\left(2\gamma-\frac{(\log 2)^{2}}{2}\right)-\log 2\log s+O(z),$
so that, using Lemma 4, we have that
$\underset{{{z}={0}}}{\textnormal{Res}}\left\{s^{-z}\Gamma(z)\eta(1+z)\zeta_{P}(z)\right\}$ $=-\frac{\log 2}{\log\beta}\log s+C_{3},$
where
$C_{3}:=-\log 2\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)+\frac{1}{\log\beta}\left(2\gamma-\frac{(\log 2)^{2}}{2}\right).$
We also need to add in the contribution from $-\log(1-e^{-s})$, which using the expansions from Section 2, is, as $s\to 0^{+}$,
$-\log(1-e^{-s})=-\log s+2\gamma+O(s).$
As in the above case of $F_{P}(e^{s})$, the contributions coming from the rest of the poles are as discussed before the statement of Lemma 4. ∎
On inspection, one notices that the dominant asymptotic terms of both $\log F_{P}(e^{-s})$ and $\log G_{P}(e^{-s})$ are precisely the dominant asymptotic terms of $\log F_{2}(e^{-s})$ and $\log G_{2}(e^{-s})$ of the functions related to the Fibonacci partitions, after substituting $\beta$ for $\varphi$. Of course, this is not so surprising, as the Fibonacci numbers $F_{k}$ are just a special case of the more general sequence $P_{k}$. The property of note, here, is regarding the associated saddle points. Since the leading order behaviour is the same, the saddle points satisfy the same leading order asymptotics. In particular, Proposition 6 and Corollary 1 hold for the saddle point $\alpha_{F_{P}}$ related to non-distinct partitions of $n$ over $P_{k}$, and so also then, do (20) and (21). Similarly, Proposition 8, Corollary 3, (22) and (23), hold for the saddle point $\alpha_{G_{P}}$ related to the partial sums of distinct partitions of $n$ over $P_{k}$. We use these results, as well as this notation, below.
@@
We first consider the case of $p_{P}(n)$, so, of $F_{P}(e^{-\alpha_{F_{p}}})$. Then, by Proposition 10, Proposition 6 and Corollary 1, the small-$\alpha_{F_{P}}$ asymptotics, or, equivalently, the large-$n$ asymptotics satisfy,
@@
$=\frac{1}{2\log\beta}\left((\log n)^{2}+2\log n\log\log\beta+(\log\log\beta)^{2}+O\left(\frac{(\log n)^{2}}{n}\right)\right)$
$\qquad\qquad+\left(\log n+\log\log\beta+O\left(\frac{\log n}{n}\right)\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)$
@@
$=\frac{(\log n)^{2}}{2\log\beta}+(\log n)\left(\frac{\log\log\beta}{\log\beta}+\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)\right)$
$\qquad\qquad+\frac{(\log\log\beta)^{2}}{2\log\beta}+\left(\log\log\beta\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)$
$\qquad\qquad\qquad+C_{2}+\psi_{4}\left(\frac{\log n}{\log\beta}\right)+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right),$
@@
$F_{P}(e^{-\alpha_{F_{P}}})=\psi_{5}\left(\frac{\log n}{\log\beta}\right)n^{\left(\frac{\log n}{2\log\beta}+\frac{\log\log\beta}{\log\beta}+\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)}\left(1+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right)\right),$
@@
$\psi_{5}\left(\frac{\log n}{\log\beta}\right):=\exp\left(\frac{(\log\log\beta)^{2}}{2\log\beta}+\left(\log\log\beta\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)+C_{2}+\psi_{4}\left(\frac{\log n}{\log\beta}\right)\right).$
@@
We now consider the number $q_{P}(n)$ of distinct partitions over $P_{k}$, and so the asymptotic behaviour of $G_{P}(e^{-\alpha_{G_{p}}})$. Then, by Proposition 10, Proposition 8 and Corollary 3, the small-$\alpha_{G_{P}}$ asymptotics, or, equivalently, the large-$n$ asymptotics satisfy,
$\log G_{P}(e^{-\alpha_{G_{p}}})$ $=(\log n)\left(\frac{\log 2}{\log\beta}+1\right)+\left(\frac{\log 2}{\log\beta}+1\right)\log\left(\frac{\log 2}{\log\beta}+1\right)$
$\qquad\qquad+C_{4}+\psi_{6}\left(\frac{\log n}{\log\beta}\right)+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right),$
where $\psi_{6}(x)$ is an explicit $1$-periodic function. Thus, we have
$G_{P}(e^{-\alpha_{G_{p}}})=\psi_{7}\left(\frac{\log n}{\log\varphi}\right)n^{\left(\frac{\log 2}{\log\beta}+1\right)}\left(1+O\left(n^{-\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}}\right)\right),$
where $\psi_{7}(x)$ is an explicit $1$-periodic function. The asymptotic for $G_{P}(e^{-\alpha})$ combined with the contributions from equations (22) and (23) finishes the proof of the theorem. ∎

### 15 Dec 2023 v2

@@ 4. Partitions over general linear recurrences with a dominant root
The results of the previous sections on Fibonacci partitions can be generalised to positive recurrence sequences $P_{n}$ with an irreducible characteristic polynomial having a positive dominant real root and such that $P_{1}=1$. We will also assume that the values $P_{n}$ are distinct. Here, we require $P_{1}=1$ so that there always exists a non-distinct partition of $n$ over $P_{n}$. The restriction on the values of $P_{n}$ being distinct is not a very strict one—since the recursion has a dominant real root there are at most finitely many repeated values, so an analysis analogous to moving between $F(x)$ and $F_{2}(x)$ is possible, and uncomplicated. For such $P_{n}$, we wish to asymptotically understand the number of solutions to
@@
As before, we let $p_{P}(n)$ denote the number of non-distinct partitions of $n$ over the sequence $P_{n}$, that is, solutions to (1) in nonnegative integers $a_{k}$.
The case of partitions over $P_{n}$ is carried out exactly as in the case with the Fibonacci numbers $F_{n}$, but with the Fibonacci zeta function replaced by the zeta function $\zeta_{P}(z)$ which is the meromorphic continuation of the Dirichlet series $\sum_{k\geqslant 1}P_{k}^{-z}$ ($\Re(z)>0$). Here, we consider the generating function,
$F_{P}(x):=\sum_{n\geqslant 0}p_{P}(n)x^{n}=\prod_{k\geqslant 1}\left(1-x^{P_{k}}\right)^{-1}.$
@@
  • a triple pole at $z=0$ in the case of $F_{P}(e^{-s})$,
@@
Now, it is clear from the previous analysis on Fibonacci partitions that the main contribution will come from the pole at $z=0$, which comes only from $\boldsymbol{k}=\boldsymbol{0}=(0,\ldots,0).$ The negative real poles will give a cumulative contribution of $O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}})$ for any fixed $\varepsilon>0$, since $s\log s=O(s^{1-\varepsilon})$ for any fixed $\varepsilon>0$ as $s\to 0^{+}$. As well, the simple poles off the real line contribute a function $f_{2}(s)$ toward $\log F_{P}(e^{-s})$, that satisfies $f_{2}(s)=f_{2}(\beta s)+O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}})$. It remains to obtain the contributions from the pole at $z=0$. For these, we require the following result.
@@
In the following result, we use Lemma 4 to determine the asymptotic behaviour of the function $\log F_{P}(e^{-s})$ as $s\to 0^{+}$. All of the expansions of the functions involved, $s^{-z}$, $\Gamma(z)$ and $\zeta(1+z)$, have been noted somewhere in the previous sections of this work—we use them below without further reference.
@@
For any fixed $\varepsilon\in(0,1)$, as $s\to 0^{+}$, the function $F_{P}(z)$ defined above satisfies,
@@
and $f_{2}(s)=f_{2}(\beta s)+O\big{(}s^{\min\{{\log(\beta/|\beta_{2}|)}/{\log\beta},1-\varepsilon\}})$.
@@
Note that near $z=0$, we have
@@
where $C_{2}:=\frac{1}{\log\beta}\left(\frac{1}{2}\left(\gamma^{2}+\frac{\pi^{2}}{6}\right)+\gamma^{2}-\gamma_{1}\right)-\left(\frac{\log\lambda}{\log\beta}+\frac{1}{2}\right)2\gamma+\left(\frac{(\log\lambda)^{2}}{2\log\beta}+\frac{\log\lambda}{2}+\frac{\log\beta}{12}+C_{1}\right).$ The contributions coming from the rest of the poles are as discussed before the statement of Lemma 4. ∎
On inspection, one notices that the dominant asymptotic terms of $\log F_{P}(e^{-s})$ are precisely the dominant asymptotic terms of $\log F_{2}(e^{-s})$, the function related to the Fibonacci partitions, after substituting $\beta$ for $\varphi$. Of course, this is not so surprising, as the Fibonacci numbers $F_{k}$ are just a special case of the more general sequence $P_{k}$. The property of note, here, is regarding the associated saddle point. Since the leading order behaviour is the same, the saddle points satisfy the same leading order asymptotics. In particular, Proposition 6 and Corollary 1 hold for the saddle point $\alpha_{F_{P}}$ related to non-distinct partitions of $n$ over $P_{k}$, and so also then, do (20) and (21). We use these results, as well as this notation, below.
@@
By Proposition 9, Proposition 6 and Corollary 1, the small-$\alpha_{F_{P}}$ asymptotics, or, equivalently, the large-$n$ asymptotics satisfy,
@@
$=\frac{1}{2\log\beta}(\log n)^{2}\left(1+O\left(\frac{1}{\log n}\right)\right)$
$\qquad\qquad+\left(\log n+\log\log\beta+O\left(\frac{1}{\log n}\right)\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)$
@@
$=\frac{(\log n)^{2}}{2\log\beta}+(\log n)\left(\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)\right)$
$\qquad\qquad+\left(\log\log\beta\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)$
$\qquad\qquad\qquad+C_{2}+\psi_{4}\left(\frac{\log n}{\log\beta}\right)+O\left(\frac{1}{\log n}\right),$
@@
$F_{P}(e^{-\alpha_{F_{P}}})=\psi_{5}\left(\frac{\log n}{\log\beta}\right)n^{\left(\frac{\log n}{2\log\beta}+\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}+O\left(\frac{1}{\log n}\right)\right)}\left(1+O\left(\frac{1}{\log n}\right)\right),$
@@
$\psi_{5}\left(\frac{\log n}{\log\beta}\right):=\exp\left(\left(\log\log\beta\right)\left(\frac{2\gamma}{\log\beta}-\frac{\log\lambda}{\log\beta}-\frac{1}{2}\right)+C_{2}+\psi_{4}\left(\frac{\log n}{\log\beta}\right)\right).$